OpenStudy (anonymous):

Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5). Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (anonymous):

First, get the slope using the slope formula:\[\frac{ y _{1} - y _{2} }{ x _{1} - x _{2} } = m\]

OpenStudy (anonymous):

So slope is 3? now what i just need the answer and part 2

OpenStudy (anonymous):

Now, take that slope, which is "m" and which you calculated in the previous post and use the point-slope formula:\[y - y _{1} = m(x - x _{1})\] It doesn't matter which point you use here at this step for (x1, y1) Yes, slope is "3". Good work!

OpenStudy (anonymous):

I got y=3x-11....but it that in standard form?

OpenStudy (anonymous):

Careful with that negative sign, You're close but you flipped the sign. And that is slope-intercept form not standard form. Slope-intercept form is y = mx + b. Standard form is ax + by = c

OpenStudy (anonymous):

So what do i do ?

OpenStudy (anonymous):

BTW, whichever point you use in step 2, you get the same line, which only makes sense.

OpenStudy (anonymous):

y - (-4) = 3[x - (-1)] -> y + 4 = 3(x + 1) -> y = 3x - 1

OpenStudy (anonymous):

That's the equation in slope-intercept form. In standard form, -3x + y = -1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

You're welcome!